# Finslerian structure of black holes

### v1

@@ III.2 Stationary black holes
From this relation, we see that the ergosurfaces correspond effectively to the critical drift case $||W||=1$. Since the solutions of (31) are $\alpha=1\pm||W||^{-1}$, we have that on the ergosurfaces we effectively always have a trapped direction, which coincides with the local direction of $W^{i}$. More precisely, the trapped propagation direction ($dx^{i}=0$) for null geodesics corresponds to the origin of $T_{\mathbf{x}}\mathcal{M}$, and the origin belongs to the indicatrix only in the critical case. However, unlike the static case, for rotating black hole this trapped direction is not the radial one.
@@
Figure 2: Depiction of the Finslerian indicatrices for the ergosphere and the outer event horizon along the $x$ axis in the equatorial plane of a Kerr black hole ($a>0$) in Natario’s coordinates (21). When $a<0$, the figure is reflected with respect to the $x$ axis, corresponding to the transformation $y\to-y$ in (III.2). In all cases, the indicatrices are horizontal ellipses, with their centers displaced according to (III.2). Notice that the centers of the indicatrices are also displaced in the azimuthal direction (vertical in the figure) due to the well-known frame-dragging effect. Right: The weak drift regime, in which null geodesics starting from a given point can propagate in all directions. Center: The critical regime corresponding to the ergosurface. The propagation of null geodesics is restricted to directions that are positively aligned with the drift, as shown in equation (5). In contrast with the Schwarzschild ($a=0$) case, the half-space of possible propagation directions (bounded by the blue dotted line) is not tangent to the ergosphere, meaning that null geodesics starting on the ergosphere can still escape to infinity. Nevertheless, they are subjected to $d\phi_{BL}\geq 0$, meaning that null geodesics are corotating with the black-hole in the usual Boyer-Lindquist coordinates. This is the only case where we can have trapped directions $(dx^{i}=0)$. Left: The event horizon. In this case, the propagation is similarly limited to the directions that are positively aligned with the drift and, moreover, constrained to lie within the conic region outlined in (7) (blue dotted line). Notice the cone is such that $dr\leq 0$ on the event horizon, meaning that no null geodesics can escape from the interior of the horizon.
@@
It is convenient to consider the azimuthal Boyer-Lindquist angle (26) to understand the distinctive properties of the ergosurfaces and their interior region. For the trapped geodesic described by (30), we have
@@
demonstrating that the ergosphere corresponds to the frontier of the “corotating” region (ergosphere) where $d\phi_{BL}$ and $\Omega$ have the same sign, i.e. $d\phi_{BL}\geq 0\,(\leq 0)$ for $\Omega>0\,(<0)$. Notice (34) explicitly shows the well-known fact that the azimuthal Boyer-Lindquist angle is well defined only in the exterior region of the event horizon, in clear contrast with Natario’s azimuthal coordinate.
@@
Finally, it is worth noticing that in the limit of small angular momentum, where only the linear terms in $a$ are considered, the spatial metric (21) simplifies to a flat metric. In this case, we recover the Zermelo navigation problem associated with the slowly rotating Kerr metric discussed in reference [11]. The Finslerian indicatrices in this scenario are qualitatively similar to those shown in Fig. 2, but with the event horizon and the ergosurface positioned arbitrarily close to each other.
IV Final remarks
@@
We have shown, through explicit and relevant physical examples, that recent mathematical findings (see [7] for a recent review) regarding the Zermelo navigation problem and its underlying Finsler geometry can provide a complete description of the causal structure of black holes, answering affirmatively the question posed in Ref. [11] about the extension of the Finsler structure to the interior of black holes. Specifically, the Kropina metric, associated with the critical Zermelo navigation problem, can fully characterize the causal properties of ergosurfaces in stationary black holes and event horizons in the spherically symmetric case. On the other hand, the Lorentz-Finsler metric, which applies to the case of a strong Zermelo problem, describes the causal structure of the interior (non-stationary) region of black holes. This complete Finslerian description offers an enhanced visual understanding of the river model of black holes, presenting the behavior of null geodesics on the ergosurfaces, horizons, and their interior regions in novel geometric terms. There are many other recent mathematical results that may have physical applications within this framework; these topics are currently under investigation.
@@
Our second and final remark concerns the analog gravity program, a robust framework that explores the emergence of effective curved spacetime geometries in non-gravitational systems [41, 42, 43]. The study of Finsler geometries can be further advanced by investigating their connections to analog spacetimes, such as those realized in hydrodynamic flows [9, 20, 44, 45], Bose-Einstein condensates [44, 45, 46], and birefringent crystals [15, 47]. As we have argued, the interplay between Spacetime Physics and Finsler geometries becomes particularly evident when the metric is recast in the Painlevé-Gullstrand form. The Painlevé-Gullstrand form not only emphasizes the causal structure of the underlying spacetime, but also naturally aligns with the Zermelo framework. In analog spacetimes, this alignment is not incidental since the Painlevé-Gullstrand form is particularly suitable for systems with a preferred frame, such as fluids or optical media. By establishing a direct connection between Finsler geometries and the causal structure of black hole interiors through explicit constructions, our work lays the foundation for deeper investigations into analog horizons and ergosurfaces, offering fresh perspectives on Finslerian concepts within the framework of analog gravity.
@@ Acknowledgments
This study was financed in part by the São Paulo Research Foundation (FAPESP, Brazil) - Process Numbers 2022/15371-3 (HR), 2022/08335-0 (MR), 2021/09293-7 (AS). The authors also acknowledge partial support from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq, Brazil), Grants 315991/2023-2 (MR) and 306785/2022-6 (AS). AS also wishes to thank Vitor Cardoso and José S. Lemos for the warm hospitality at the Center for Astrophysics and Gravitation of the University of Lisbon, where part of this work was done.

### v2

@@ III.2 Stationary black holes
From this relation, we see that the ergosurfaces correspond effectively to the critical drift case $||W||=1$. Since the solutions of (39) are $\alpha=1\pm||W||^{-1}$, we have that on the ergosurfaces we effectively always have a blocked direction, i.e., a direction such that $dx^{i}=0$ for null geodesics. More precisely, the blocked propagation direction for null geodesics corresponds to the origin of $T_{\mathbf{x}}\mathcal{M}$, and the origin belongs to the indicatrix only in the critical case. However, unlike the static case, for rotating black hole this blocked direction is not the radial one.
@@
Figure 2: Depiction of the Finslerian indicatrices for the ergosphere and the outer event horizon along the $x$ axis in the equatorial plane of a Kerr black hole ($a>0$) in Natario’s coordinates (29). When $a<0$, the figure is reflected with respect to the $x$ axis, corresponding to the transformation $y\to-y$ in (III.2). In all cases, the indicatrices are horizontal ellipses, with their centers displaced according to (III.2). Notice that the centers of the indicatrices are also displaced in the azimuthal direction (vertical in the figure) due to the well-known frame-dragging effect. Right: The weak drift regime, in which null geodesics starting from a given point can propagate in all directions. Center: The critical regime corresponding to the ergosurface. The propagation of null geodesics is restricted to directions that are positively aligned with the drift, as shown in equation (5). In contrast with the Schwarzschild ($a=0$) case, the half-space of possible propagation directions (bounded by the blue dotted line) is not tangent to the ergosphere, meaning that null geodesics starting on the ergosphere can still escape to infinity. Nevertheless, they are subjected to $d\phi_{BL}\geq 0$, meaning that null geodesics are corotating with the black-hole in the usual Boyer-Lindquist coordinates. This is the only case where we can have blocked directions $(dx^{i}=0)$. Left: The event horizon. In this case, the propagation is similarly limited to the directions that are positively aligned with the drift and, moreover, constrained to lie within the conic region outlined in (7) (blue dotted line). Notice the cone is such that $dr\leq 0$ on the event horizon, meaning that no null geodesics can escape from the interior of the horizon. This is a concrete realization, for a Kerr black hole, of the wind-Finslerian description of the causal structure of spacetimes introduced in [7].
@@
It is convenient to consider the azimuthal Boyer-Lindquist angle (34) to understand the distinctive properties of the ergosurfaces and their interior region. For the blocked geodesic described by (38), we have
@@
demonstrating that the ergosurface corresponds to the frontier of the “corotating” region (ergosphere) where $d\phi_{BL}$ and $\Omega$ have the same sign, i.e. $d\phi_{BL}\geq 0\,(\leq 0)$ for $\Omega>0\,(<0)$. Notice (42) explicitly shows the well-known fact that the azimuthal Boyer-Lindquist angle is well defined only in the exterior region of the event horizon, in clear contrast with Natario’s azimuthal coordinate.
@@
The limit of small angular momentum is particularly interesting from a physical perspective. In this limit, where only linear terms in $a$ are considered, the spatial metric (29) simplifies to a flat metric. As a result, we recover the Zermelo navigation problem associated with the slowly rotating Kerr metric, as discussed in reference [12]. In this scenario, the Finslerian indicatrices are qualitatively similar to those shown in Fig. 2, with the event horizon and the ergosurface positioned very close to each other.
Finally, it is noteworthy that the Finslerian indicatrices depicted in Fig. 2, particularly in the critical and strong drift cases, suggest potential interpretations of superradiance and the Penrose process through the Finslerian framework. In simple terms, these intimately connected phenomena [45] allow for the extraction of energy from Kerr black holes. Superradiance is a scattering phenomenon in which outgoing waves carry more energy than the incoming ones. Similarly, the Penrose process typically involves particle collisions near the horizon, where the black hole absorbs a particle with negative energy, allowing a more energetic particle to escape to infinity The existence of the ergosphere, where no observer can remain static, is crucial for both phenomena. We have established that the ergosurface $r=2M$, which forms the external boundary of the ergosphere, corresponds to the critical drift region in Natario’s coordinates.
Nevertheless, it is important to note that superradiance can also occur in flat spacetimes, even in the absence of horizons or an ergosurface [45, 46]. A particularly illustrative example is the Vavilov-Cherenkov effect, in which radiation is emitted by a charge moving faster than the speed of light in a dielectric medium. This phenomenon, which has a long and rich history [47, 48], can be explained in terms of spontaneous superradiance involving the formation of a Mach-like cone that characterizes a shock wave [45]. The shock waves associated with superluminal motion in a dielectric can be understood as a superradiant amplification of electromagnetic waves [49], and similar effects can also be observed with sound waves and gravitational interactions [50, 51]. In all these cases, the Mach-like cone has an opening angle
$\vartheta=\pi-2\cos^{-1}\left(\frac{v_{ph}}{v}\right)=2\cos^{-1}\sqrt{1-\left(\frac{v_{ph}}{v}\right)^{2}},$ (43)
where $v_{ph}$ denotes the phase velocity of the radiation in the external medium and $v=||\mathbf{V}||$, with $\mathbf{V}$ being the particle velocity.
The key observation here is that the Vavilov-Cherenkov effect can also be qualitatively understood from the comoving reference frame, where the particle is at rest and the medium is moving with velocity $-\mathbf{V}$. We can follow [52] and describe the Vavilov-Cherenkov radiation in terms of plane waves moving in a medium with phase velocity $v_{ph}$
$\varphi_{\mathbf{k}}(t,\mathbf{r})=e^{-i\left(v_{ph}kt-\mathbf{k}\cdot\mathbf{r}\right)},$ (44)
where $\mathbf{k}$ stands for the wave vector and $k=||\mathbf{k}||$. From the comoving reference frame, such modes are perceived as harmonic excitations $\varphi_{\mathbf{k}}(t(\tau),\mathbf{r}(\tau))=e^{-i\omega\tau}$ with frequencies
$\omega=\gamma\left(v_{ph}k-\mathbf{V}\cdot\mathbf{k}\right),$ (45)
where $\tau$ is the comoving proper time and $\lambda$ is the Lorentz factor. The Vavilov-Cherenkov effect is associated with the negative frequencies inside the Vavilov-Cherenkov cone, see [52] for further details. It turns out that the negative frequencies $\omega$ in (45) occur exactly inside the cone defined by (43). It is easy to see that, for the cases with $v>v_{ph}$, in the comoving frame, the modes emitted inside the cone (43) have negative frequency (energy), while those emitted outside the cone have positive frequency. Superradiant excitations are known to arise from the energy balance involving positive and negative frequency modes [52]. However, from the comoving reference, light rays in the external moving medium will behave locally exactly as predicted by the wind-Finslerian indicatrices with $\mathbf{W}=-\frac{1}{v_{ph}}\mathbf{V}$. In other words, once the critical drift region is crossed and the directions of the null geodesics become constrained by the cone structures described in equations (5) and (7), the situation closely resembles that of spontaneous superradiance in flat spacetimes. From (7), with the standard definition of the angle between two vectors in terms of their scalar product, the opening angle of the Finsler cones is given by
$\vartheta=2\cos^{-1}\sqrt{1-\frac{1}{||W||^{2}}},$ (46)
which coincides with (43) in the present case. As expected, the cone forms at the critical region where $||W||=1$, and becomes more acute as we plunge into the strong drift region $||W||>1$. For static black holes, the critical region coincides with the horizon and, as a result, any potential superradiance arising from negative energy radiation emitted within the cone will remain confined to the interior of the black hole. In contrast, for stationary spacetimes such as the Kerr black hole spacetime, the critical region corresponds to the ergosurface, which lies outside the event horizon. This creates a region, the ergosphere, from which radiation emitted within the cone can escape to infinity (see Fig. 2), potentially leading to superradiance. This scenario paves the way for a new and interesting interpretation of superradiance in terms of the wind-Finslerian structure of black holes.
IV Analog models
Analog gravity is a robust framework that explores the emergence of effective curved spacetime geometries in non-gravitational systems [53, 54, 55, 56]. These analog spacetimes are particularly suitable for analysis from a wind-Finslerian perspective, as they often involve effective metrics of the Painlevé-Gullstrand form (III). For many analog spacetimes, such as those based on hydrodynamic flows and Bose-Einstein condensates, the “river model” description is not merely an alternative interpretation, but rather the genuine underlying physics of the system.
An especially illustrative class of analog models where the Finslerian formalism finds direct and novel application is that of vortex flows in fluids, often employed as analogs of rotating black holes. This is the case, for instance, of the configuration known as the “draining bathtub” model, described in the standard Euclidean two-dimensional space by the background velocity field
${W}=-\frac{A}{r}\hat{r}+\frac{B}{r}\hat{\varphi},$ (47)
where the radial component $A>0$ controls the inward drift and $B$ encodes the vortex circulation. The associated acoustic metric, derived via linear perturbations of the velocity potential, has a Painlevé–Gullstrand form (III), see [57] for details. In this case,
$||W||^{2}=\frac{A^{2}+B^{2}}{r^{2}},$ (48)
and the Finslerian indicatrix for null geodesics starting along the $x$ axis reads
$\left(dx+{\frac{A}{r}}\right)^{2}+\left(dy-{\frac{B}{r}}\right)^{2}=1.$ (49)
The overall behavior of this indicatrix is qualitatively similar to those of the previous section. The strong drift regime emerges where the inward radial flow exceeds the local wave propagation speed $c_{w}=1$, resulting in a trapping horizon. This occurs at $r_{h}=A$, where the underlying geometry is governed by a Lorentz-Finsler metric. On the other hand, an analog ergosurface is found at the critical region $r_{e}=\sqrt{A^{2}+B^{2}}$, where $||W||=1$. Thus, the presence of the azimuthal component of the drift leads to the formation of an ergoregion $r_{h}<r<r_{e}$, analogous to that of a Kerr black hole, where the co-rotating frame speed exceeds the local wave speed in the azimuthal direction.
This analog ergoregion possesses the necessary geometric features for superradiant scattering, and, indeed, an active area of research with experimental results [58] has developed around this topic. Whenever an ergoregion is present together with a dissipation mechanism, wave modes with azimuthal number $m$ and frequency $\omega$ satisfying $\omega<m\Omega$, where $\Omega$ is the local angular velocity of the flow, are amplified upon reflection due to the presence of negative-energy modes in the comoving frame [53, 57, 54]. The Finsler indicatrix (49), being displaced and skewed by the azimuthal drift as in Fig. 2, encodes this kinematic conditions geometrically, reflecting the frame-dragging and the energy balance that underpins superradiance. In this way, Finsler geometry offers a covariant and geometrically transparent description of the interplay between ergoregion formation and superradiance phenomena in vortex-based analog black hole systems. Furthermore, this framework may be instrumental in analyzing stability properties and backreaction effects in experiments involving dispersive corrections.
We finish by commenting that flows of the type (47), with $A<0$ and $B=0$, correspond to the standard case of the circular hydraulic jump, a canonical example of an analog white hole horizon. Hydraulic jumps form when the propagation of surface waves on a thin liquid layer undergoes a transition from a supercritical to a subcritical regime, marking a causal boundary for counterstream wave propagation [59]. In the shallow-water limit, the dispersion relation for gravity surface waves reduces to a relativistic form [60, 61]
$(\omega-\mathbf{W}\cdot\mathbf{k})^{2}\approx||\mathbf{k}||^{2}=k^{2},$ (50)
where $\mathbf{W}$ is the local background flow velocity, and $\omega$ and $\mathbf{k}$ denote the wave frequency and the wave vector, respectively. It turns out that the dispersion relation (50) admits an unexpected but clear interpretation in Finslerian terms. Notice that a wave traveling in two dimensions, with frequency $\omega$ and wave vector $\mathbf{k}$, moves along the vector $\mathbf{V}=dx\hat{\imath}+dy\hat{\jmath}=\frac{\omega}{k^{2}}\mathbf{k}$. In terms of the vector $\mathbf{V}$, for a radial flow $\mathbf{W}=W\hat{r}$, with $W=\frac{|A|}{r}>0$, we find
$\left(dx-W\right)^{2}+dy^{2}=1$ (51)
for waves starting on the $x$ axis. Thus, the dispersion relation (50) defines an indicatrix, exhibiting the same qualitative behavior discussed in the previous sections. The critical region $r_{h}=|A|$ acts as a white hole horizon and determines the location of the hydraulic jump of the flow [59]. Outside the hydraulic jump, we are in the weak drift regime, where surface waves can propagate in all directions. At the radius of the hydraulic jump $r=R_{j}$, the drift matches the wave speed so that only radially outward propagation is permitted. Finally, in the strong drift regime, corresponding to the interior of the hydraulic jump, only outgoing geodesics exist, mimicking the causal disconnection characteristic of white hole interiors. As one can see, the wind-Finslerian picture offers a rigorous and geometrically transparent framework for describing horizon formation, causal disconnection, and wave dynamics also for analog models involving hydraulic jumps.
V Final remarks
@@
We have shown, through explicit and relevant physical examples, that recent mathematical findings (see [7] for the main results and a comprehensive review) regarding the Zermelo navigation problem and its underlying wind-Finslerian structure can provide a complete description of the causal structure of black holes, answering affirmatively the question posed in Ref. [12] about the extension of the Finsler structure to the interior of black holes. Such results are a direct consequence of the wind-Finslerian structures introduced and developed in [7]. Specifically, the Kropina metric, associated with the critical Zermelo navigation problem, can fully characterize the causal properties of ergosurfaces in stationary black holes and event horizons in the spherically symmetric case. On the other hand, the Lorentz-Finsler metric, which applies to the case of a strong Zermelo problem, describes the causal structure of the interior (non-stationary) region of black holes. This complete wind-Finslerian description offers an enhanced visual understanding of the river model of black holes, presenting the behavior of null geodesics on the ergosurfaces, horizons, and their interior regions in novel geometric terms. There are many other recent mathematical results that may have physical applications within this framework; these topics are currently under investigation.
@@
Our second and final remark concerns the analog gravity program. Besides the issues discussed in Section IV, the study of Finsler geometries can be further advanced by investigating their connections to analog spacetimes, such as those realized in hydrodynamic flows [10, 22, 62, 63], Bose-Einstein condensates [62, 63, 64], and birefringent crystals [17, 65]. As we have argued, the interplay between Spacetime Physics and Finsler geometries becomes particularly evident when the metric is recast in the Painlevé-Gullstrand form. The Painlevé-Gullstrand form not only emphasizes the causal structure of the underlying spacetime, but also naturally aligns with the wind-Finslerian framework. In analog spacetimes, this alignment is not incidental since the Painlevé-Gullstrand form is particularly suitable for systems with a preferred frame, such as fluids or optical media. By establishing a direct connection between Finsler geometries and the causal structure of black hole interiors through explicit constructions, our work lays the foundation for deeper investigations into analog horizons and ergosurfaces, offering fresh perspectives on wind-Finslerian concepts within the framework of analog gravity.
@@ Acknowledgments
This study was financed in part by the São Paulo Research Foundation (FAPESP, Brazil) - Process Numbers 2022/15371-3 (HR), 2022/08335-0 (MR), 2021/09293-7 (AS). The authors also acknowledge partial support from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq, Brazil), Grants 315991/2023-2 (MR) and 306785/2022-6 (AS). AS also wishes to thank Vitor Cardoso and José S. Lemos for the warm hospitality at the Center for Astrophysics and Gravitation of the University of Lisbon, where part of this work was done, and S. Carneiro for helpful discussions.